# Théodor Lemerle, Nicolas Obin, Axel Roebel

### v1

@@ Abstract
Recent advancements in text-to-speech (TTS) powered by language models have showcased remarkable capabilities in achieving naturalness and zero-shot voice cloning. Notably, the decoder-only transformer is the prominent architecture in this domain. However, transformers face challenges stemming from their quadratic complexity in sequence length, impeding training on lengthy sequences and resource-constrained hardware. Moreover they lack specific inductive bias with regards to the monotonic nature of TTS alignments. In response, we propose to replace transformers with emerging recurrent architectures and introduce specialized cross-attention mechanisms for reducing repeating and skipping issues. Consequently our architecture can be efficiently trained on long samples and achieve state-of-the-art zero-shot voice cloning against baselines of comparable size.
@@ 1.2 Linear surrogate of decoder transformer
(a) In decoder-only LM TTS models [7, 9, 15], attention scores either boil down to cross-attention or local reasoning so that a large proportion of tokens is never attended.
@@ 1.3 TTS as conditional codec language modeling
$p(\mathbf{y}|\mathbf{x})=\prod_{t=0}^{T}\prod_{q=1}^{Q}p(\mathbf{y_{q,t}}|\mathbf{x},\mathbf{y_{<q,\leq t}}).\vspace{-0.75cm}$ (3)
@@ 2 Small-E
This section presents Small-E [^1] A demo page with speech samples is available at: https://theodorblackbird.github.io/blog/demo_alpha/, a multi-speaker neural TTS with zero-shot voice cloning capabilities. Small-E belong to the family of neural codec language model such as [7, 9, 15, 10]. In contrast with previous TTS codec LM model that leverages decoder-only (GPT) transformers, Small-E relies on encoder-decoder architecture. Indeed, we observed that previous decoder-only transformers tend to behave internally as an encoder-decoder (see Figure 1) leading to a potential waste of compute. The general architecture is presented in Figure 2, text is encoded through a non-causal transformer, audio is encoded with a stack of LCLM blocks. Both encoders outputs are fed to the cross-attention that learns to align text to audio. The audio decoder (same as the audio encoder) takes audio embeddings and cross-attention output. The decoder output is projected to logits.
@@ 2.1 Model architecture
It is important to note that the positional embedding $\mathbf{P}$ is not superposed onto any latent vector but rather materialized independently, constraining the model to accurately encode positional information to effectively attend to the text content. We set $d_{b}$ to be significantly smaller than the model dimension ($d_{b}\leq 64$) to keep additional operations negligible. This approach is reminiscent of Location Sensitive Attention [27], while being an order of magnitude faster due to using only parallelizable operations. The optimization objective is cross-entropy loss between original RVQ codec and logits prediction.
@@ 3.1 Dataset
Small-E was trained on Librilight medium [29], consisting of approximately 5,000 hours of multi-speaker English speech recordings reading audio books, collected from LibriVox. We used the provided recipes to get samples of approximately 25 seconds. Speech utterances were transcribed textually using Ocotillo [30] speech recognition system. Validation set is made of 2,000 random utterances. For testing, the proposed model and the benchmark models (see below) were compared on the LibriTTS test split [29], in particular by insuring training and testing speakers do not overlap.
@@ 4 Results and Discussion
Model Audio token per second (kT/s) $\uparrow$ ppl $\downarrow$
@@
Table 2: Position-Aware Cross-Attention impact on 100 utterances. Number of utterances that contains at least one skip or one repetition.

### v2

@@ Abstract
Recent advancements in text-to-speech (TTS) powered by language models have showcased remarkable capabilities in achieving naturalness and zero-shot voice cloning. Notably, the decoder-only transformer is the prominent architecture in this domain. However, transformers face challenges stemming from their quadratic complexity in sequence length, impeding training on lengthy sequences and resource-constrained hardware. Moreover they lack specific inductive bias with regards to the monotonic nature of TTS alignments. In response, we propose to replace transformers with emerging recurrent architectures and introduce specialized cross-attention mechanisms for reducing repeating and skipping issues. Consequently our architecture can be efficiently trained on long samples and achieve state-of-the-art zero-shot voice cloning against baselines of comparable size. Our implementation and demos are available at https://github.com/theodorblackbird/lina-speech.
@@ 1.2 Linear surrogate of decoder transformer
(a) In decoder-only LM TTS models [7, 9, 15], attention scores either boil down to cross-attention or local reasoning so that a large proportion of tokens are not attended.
@@ 1.3 TTS as conditional codec language modeling
$p(\mathbf{y}|\mathbf{x})=\prod_{t=0}^{T}\prod_{q=1}^{Q}p(\mathbf{y_{q,t}}|\mathbf{x},\mathbf{y_{<q,\leq t}}).$ (3)
@@ 2 Small-E
This section presents Small-E, a multi-speaker neural TTS with zero-shot voice cloning capabilities. Small-E belong to the family of neural codec language model such as [7, 9, 15, 10]. In contrast with previous TTS codec LM model that leverages decoder-only (GPT) transformers, Small-E relies on encoder-decoder architecture. Indeed, we observed that previous decoder-only transformers tend to behave internally as an encoder-decoder (see Figure 1) leading to a potential waste of compute. The general architecture is presented in Figure 2, text is encoded through a non-causal transformer, audio is encoded with a stack of LCLM blocks. Both encoders outputs are fed to the cross-attention that learns to align text to audio. The audio decoder (same as the audio encoder) takes audio embeddings and cross-attention output. The decoder output is projected to logits.
@@ 2.1 Model architecture
It is important to note that the positional embedding $\mathbf{P}$ is not superposed onto any latent vector but rather materialized independently, constraining the model to accurately encode positional information to effectively attend to the text content. We set $d_{b}$ to be significantly smaller than the model dimension ($d_{b}\leq 64$) to keep additional operations negligible. This approach is reminiscent of Location Sensitive Attention [27], while being an order of magnitude faster due to the efficiency of LCLM blocks. The optimization objective is cross-entropy loss between original RVQ codec and logits prediction.
@@ 3.1 Dataset
Small-E was trained on Librilight medium [29], consisting of approximately 5,000 hours of multi-speaker English speech recordings reading audio books, collected from LibriVox. We used the provided recipes to get samples of approximately 25 seconds. Speech utterances were transcribed textually using Ocotillo [30] speech recognition system. Validation set is made of 2,000 random utterances. For evaluation, the proposed model and the benchmark models (see below) were compared on the LibriTTS test split [29], in particular by insuring training and testing speakers do not overlap.
@@ 4 Results and Discussion
Model Audio token per second (kT/s)  [^1] Using EnCodec[31] at 3kbps with delaying scheme[25], it consists of 4 tokens generated in parallel per decoding step. $\uparrow$ ppl $\downarrow$
@@
Table 2: Position-Aware Cross-Attention impact on 100 utterances. Number of utterances that contains at least one skip/repetition.